# Reconstructing Subnational Labor Indicators in Colombia: An Integrated Machine and Deep Learning Approach

### v1

@@ 2.5 Accounting Identities in Predictive Labor Modeling
Source: Author’s construction.
@@ 3.2.1 Adjustment of World Bank Labor Estimates
However, global estimation frameworks often fail to capture dynamics specific to Colombia. For example, the surge in unemployment following the 1998–1999 recession, the mild impact of the 2008-2009 financial crisis, and the sharp shock of COVID-19 in 2020 are only partially reflected Medina et al., (2013); Grosh et al., (2014); Alvarez & Pizzinelli, (2021); DANE, (2025); BanRep, (2025). To address these deficiencies, we structurally recalibrate the series: smoothing employment and unemployment to align with macroeconomic cycles, reconstructing participation based on its empirical relationship with employment, and deriving aggregates using official projections of the working-age population.
@@ 4.1.1 Main Labor Market Series
Source: Author’s calculations.
@@
Source: Author’s calculations.
@@ 4.1.2 Informality Estimation Results
informality rate 0.0149 0.0111 2.0045
Source: Author’s calculations.
Table 5: In sample RMSE, MAE y MAPE (%) per main city and metropolitan areas (2007-2024).
@@
Armenia 0.0145 0.0112 2.3244
Barranquilla A.M. 0.0189 0.0145 2.5822
Bogotá D.C. 0.0143 0.0118 3.1117
Bucaramanga A.M. 0.0107 0.0084 1.6821
Cali A.M. 0.0109 0.0087 1.7021
Cartagena 0.0161 0.0116 1.9597
Cúcuta A.M. 0.0163 0.0128 2.0054
Florencia 0.0134 0.0106 1.6559
Ibagué 0.0128 0.0100 1.7655
Manizales A.M. 0.0114 0.0086 2.2429
Medellín A.M. 0.0121 0.0097 2.2467
Montería 0.0144 0.0116 1.7431
Neiva 0.0157 0.0126 2.2775
Pasto 0.0139 0.0111 1.8029
Pereira A.M. 0.0125 0.0091 1.9799
Popayán 0.0167 0.0124 2.0179
Quibdó 0.0196 0.0139 2.1699
Riohacha 0.0168 0.0117 1.8108
Santa Marta 0.0092 0.0070 1.0404
Sincelejo 0.0212 0.0133 1.8837
Tunja 0.0159 0.0110 2.4118
Valledupar 0.0166 0.0126 1.8879
Villavicencio 0.0134 0.0108 1.7992
Source: Author’s calculations.
@@ 4.1.3 Synthesis of Validation Results
Table 2 reports the validation metrics for the six core labor variables reconstructed through temporal disaggregation and multi-output regression. For these variables, the Mean Absolute Percentage Error (MAPE) remains below 2.3%, with the lowest relative errors observed for the Participation in the Economically Active Population (1.2%) and employed (1.3%).The aggregate error across all variables results in a MAPE of 1.8%.
@@ 4.2 National Labor Trends (1993-2025)
Source: Author’s calculations.
@@
Source: Author’s calculations.
@@
Source: Author’s calculations.
@@
Source: Author’s calculations.
@@
The 2008 global recession resulted in only mild adjustments. Relative to the years 2006-2007, participation and employment increased slightly, unemployment remained stable, and inactivity decreased marginally. Informality also decreased, in line with favorable external conditions and counter-cyclical policies. No evidence of long-term scarring has been observed.
The COVID-19 pandemic resulted in an abrupt and synchronized economic collapse. In comparison to 2018-2019, participation and employment declined sharply, unemployment surged and inactivity increased. Informal employment rose following a decade of decline, highlighting the vulnerability of urban service-sector jobs to lockdowns and the limited capacity for telework. This episode combined immediate disruption with regressive structural effects.
@@
Source: Author’s calculations.
@@
Source: Author’s calculations. Pairwise correlations between indicators were computed using Spearman’s rank correlation coefficient, which is robust to non-linear monotonic relationships and less sensitive to outliers.
@@
Source: Author’s calculations. Granger causality tests were conducted allowing up to six lags, with the optimal lag for each variable pair selected via the Bayesian Information Criterion (BIC).
@@ 4.3.1 Historical Analysis of Departmental Labor Market Indicators
Source: Author’s estimations.
@@
Source: Author’s estimations.
@@
Source: Author’s estimations.
@@
Source: Author’s estimations.
@@
Source: Author’s estimations.
@@ 4.3.2 Labor Market Profiles with a Quality Index
Source: Author’s estimations.
@@
Source: Author’s estimations.
@@ Acknowledgments
All reconstructed labor market indicators produced in this study, including monthly national and departmental series from 1993 to 2025, are intended to be made publicly available through a GitHub repository archived on Zenodo. The dataset will include machine-readable files, methodological documentation, and version-controlled metadata. A persistent DOI will be assigned upon public release, and subsequent updates will adhere to semantic versioning. All input data used in the reconstruction process, including GEIH labor annexes, World Bank indicators, and ILOSTAT aggregates, are publicly accessible through their respective institutional repositories.
@@ Appendix B.8 Validation Outputs
Note: RMSE values are computed by comparing annualized disaggregated series to GEIH national estimates.
Source: Author’s calculations based on harmonized World Bank-ILO estimates and official GEIH series from DANE.
@@ Appendix C.4 Temporal Disaggregation of Annual Signals
For cases where only annual city totals $v_{c,t}$ are observed, Chow-Lin temporal disaggregation with AR(1) errors:
@@ Appendix D.2 Temporal Disaggregation of Annual Departmental Indicators
We employ the Chow-Lin method, utilizing average conversion and an AR(1) error structure:
@@ Appendix E.1 Macroeconomic Variables
  • TRM (Tasa Representativa del Mercado): Monthly average of the representative exchange rate between the Colombian peso and the US dollar.
  • IPC (Índice de Precios al Consumidor): Headline monthly inflation index and its 12-month variation.
  • IPP (Índice de Precios al Productor): Producer price index, capturing upstream cost dynamics.
  • X (Exportaciones FOB): Monthly value of total Colombian exports (in USD).
  • M (Importaciones CIF): Monthly value of total Colombian imports (in USD).
  • UVR (Unidad de Valor Real): Monthly value of the UVR, indexed to inflation expectations and relevant for financial contracts.
  • IDXPI (Índice de Producción Industrial): Monthly industrial production index, published by DANE as a proxy for real sector activity.
All series were processed to ensure temporal continuity. Missing months were reconstructed primarily using spline-based methods (Akima/PCHIP). Simple forward fill was restricted to gaps of $\leq$2 months to avoid bias and step artifacts. Consistency was validated against official annual summaries.
@@ Appendix E.2 Institutional Wage Components
  • Minimum Legal Monthly Wage (minimum_wage): Nominal statutory minimum wage in force each January, extended uniformly across months.
  • Transport Subsidy (transport_subsidy): Monthly subsidy value granted to workers earning up to two minimum wages, updated annually.
@@ Appendix H.1.5 Validation Protocols
Appendix J.3 Official Departaments Main Cities labor market indicators
Table J.7: Subnational Labor Indicators: Departments and Cities
@@
Departmental Employment Annual 23 departments (2005-2024)
Departmental Unemployment Annual 23 departments (2005-2024)
Departmental PEA Annual 23 departments (2005-2024)
Departmental Inactivity Annual 23 departments (2005-2024)
City Employment (GEIH Annex) Monthly 32 cities + Cundinamarca (2021-2024)
City Unemployment (GEIH Annex) Monthly 32 cities + Cundinamarca (2021-2024)
City Labor Force (PEA) Monthly 32 cities + Cundinamarca (2021-2024)
City Inactivity Monthly 32 cities + Cundinamarca (2021-2024)
City Participation / Employment / Unemployment / Inactivity Rates Monthly 32 cities + Cundinamarca (2021-2024)
@@
Consumer Price Index (CPI) Monthly National (1991-2025)
Producer Price Index (PPI) Monthly National (1991-2025)
Representative Market Exchange Rate (TRM) Daily / Monthly National (1991-2025)
Monthly Inflation (variation of CPI) Monthly National (1991-2025)

### v2

@@ 2.5 Accounting Identities in Predictive Labor Modeling
Note: Author’s own construction.
@@ 3.2.1 Adjustment of World Bank Labor Estimates
However, global estimation frameworks often fail to capture dynamics specific to Colombia. For example, the surge in unemployment following the 1998–1999 recession, the mild impact of the 2008-2009 financial crisis, and the sharp shock of COVID-19 in 2020 are only partially reflected Medina et al., (2013); Grosh et al., (2014); Alvarez & Pizzinelli, (2021); DANE, (2025); BanRep, (2025). To address these deficiencies, we structurally recalibrate the series by smoothing employment and unemployment to align with macroeconomic cycles, reconstructing participation based on its empirical relationship with employment, and deriving aggregates using official projections of the working-age population.
@@ 4.1.1 Main Labor Market Series
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@ 4.1.2 Informality Estimation Results
Informality Rate (%) 0.0149 0.0111 2.0045
Note: Author’s calculations.
Table 5: In-sample RMSE, MAE and MAPE (%) per main city and metropolitan areas (2007-2024).
@@
Armenia 0.015 0.011 2.32
Barranquilla A.M. 0.019 0.015 2.58
Bogotá D.C. 0.014 0.012 3.11
Bucaramanga A.M. 0.011 0.008 1.68
Cali A.M. 0.011 0.009 1.70
Cartagena 0.016 0.012 1.96
Cúcuta A.M. 0.016 0.013 2.01
Florencia 0.013 0.011 1.66
Ibagué 0.013 0.010 1.77
Manizales A.M. 0.011 0.009 2.24
Medellín A.M. 0.012 0.010 2.25
Montería 0.014 0.012 1.74
Neiva 0.016 0.013 2.28
Pasto 0.014 0.011 1.80
Pereira A.M. 0.013 0.009 1.98
Popayán 0.017 0.012 2.02
Quibdó 0.020 0.014 2.17
Riohacha 0.017 0.012 1.81
Santa Marta 0.009 0.007 1.04
Sincelejo 0.021 0.013 1.88
Tunja 0.016 0.011 2.41
Valledupar 0.017 0.013 1.89
Villavicencio 0.013 0.011 1.80
Note: Author’s calculations.
@@ 4.1.3 Synthesis of Validation Results
Table 2 reports the validation metrics for the six core labor variables reconstructed through temporal disaggregation and multi-output regression. For these variables, the Mean Absolute Percentage Error (MAPE) remains below 2.3%, with the lowest relative errors observed for Participation in the Economically Active Population (1.2%) and employment (1.3%).The aggregate error across all variables results in a MAPE of 1.8%.
@@ 4.2 National Labor Trends (1993-2025)
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
The 2008 global recession resulted in only mild adjustments. In comparison to the years 2006-2007, participation and employment increased slightly, unemployment remained stable, and inactivity decreased marginally. Informality also decreased, consistent with favorable external conditions and counter-cyclical policies. No evidence of long-term scarring has been observed.
The COVID-19 pandemic resulted in an abrupt and synchronized economic collapse. In comparison to the years 2018-2019, participation and employment declined sharply, unemployment surged, and inactivity increased. Informal employment rose after a decade of decline, highlighting the vulnerability of urban service-sector jobs to lockdowns and the limited capacity for telework. This episode combined immediate disruption with regressive structural effects.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations. Pairwise correlations between indicators were computed using Spearman’s rank correlation coefficient, which is robust to non-linear monotonic relationships and less sensitive to outliers.
@@
Note: Author’s calculations. Granger causality tests were conducted allowing up to six lags, with the optimal lag for each variable pair selected via the Bayesian Information Criterion (BIC).
@@ 4.3.1 Historical Analysis of Departmental Labor Market Indicators
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@ 4.3.2 Labor Market Profiles with a Quality Index
Note: Author’s calculations.
@@
Note: Author’s calculations.
@@ Acknowledgments
All reconstructed labor market indicators generated in this study, including monthly national and departmental series from 1993 to 2025, are publicly available at: https://doi.org/10.5281/zenodo.16899686
The repository contains machine-readable data files, methodological documentation, and version-controlled metadata. All input data used in the reconstruction process, including GEIH labor annexes, World Bank indicators, and ILOSTAT aggregates, are publicly accessible through their respective institutional repositories.
Code Availability Statement
All source code utilized for data processing, reconstruction, and visualization is publicly available in the GitHub repository archived on Zenodo: https://doi.org/10.5281/zenodo.16899686 (also accessible at https://github.com/jaimevera1107/colombian-labor-market)
The codebase is version-controlled, thoroughly documented, and comprises reproducible scripts necessary for replicating all analyses and figures presented in this paper.
@@ Appendix B.8 Validation Outputs
Note: Author’s calculations based on harmonized World Bank-ILO estimates and official GEIH series from DANE. RMSE values are computed by comparing annualized disaggregated series to GEIH national estimates.
@@ Appendix C.4 Temporal Disaggregation of Annual Signals
For cases where only annual city totals $v_{c,t}$ are observed, Chow-Lin temporal disaggregation with AR(1) errors is employed:
@@ Appendix D.2 Temporal Disaggregation of Annual Departmental Indicators
We employ the Chow-Lin method, utilizing average conversion and an autoregressive (AR(1)) error structure:
@@ Appendix E.1 Macroeconomic Variables
  • TRM (TRM, Tasa Representativa del Mercado), Monthly average of the representative exchange rate between the Colombian peso and the US dollar.
  • CPI (IPC, Índice de Precios al Consumidor), Headline monthly inflation index and its 12-month variation.
  • PPI (IPP, Índice de Precios al Productor), Producer price index, capturing upstream cost dynamics.
  • X (FOB, Exportaciones), Monthly value of total Colombian exports (in USD).
  • M (CIF, Importaciones), Monthly value of total Colombian imports (in USD).
  • UVR (UVR, Unidad de Valor Real), Monthly value of the UVR, indexed to inflation expectations and relevant for financial contracts.
  • IPI (IPI, Índice de Producción Industrial), Monthly industrial production index, published by DANE as a proxy for real sector activity.
All series were processed to ensure temporal continuity. Missing months were reconstructed primarily using spline-based methods (Akima/PCHIP). Simple forward fill was restricted to gaps of $\leq$12 months to avoid bias and step artifacts. Consistency was validated against official annual summaries.
@@ Appendix E.2 Institutional Wage Components
  • Minimum Legal Monthly Wage (SMMLV, Salario Mínimo Mensual Legal Vigente), Annual Nominal statutory minimum wage in force each January, extended uniformly across months.
  • Transport Subsidy (ST, Subsidio de Transporte), Annual subsidy value granted to workers earning up to two minimum wages, updated annually.
@@ Appendix H.1.5 Validation Protocols
Note: Author’s own construction.
Appendix J.3 Official Departments Labor Market Indicators
Table J.7: Subnational Labor Indicators: Departments
@@
Employment Annual 33 departments (2005–2024)
Unemployment Annual 33 departments (2005–2024)
Economically Active Population (PEA) Annual 33 departments (2005–2024)
Working-Age Population (PET) Annual 33 departments (2005–2024)
Total Population Annual 33 departments (2005–2024)
Inactive Population Annual 33 departments (2005–2024)
Participation / Employment / Unemployment / Inactivity Rates Annual 33 departments (2005–2024)
Note: Author’s own construction.
Appendix J.4 Official Main Cities Labor Market Indicators
Table J.8: Subnational Labor Indicators: Main Cities
Indicator Frequency Coverage
Employment Monthly 32 cities + Cundinamarca (2021–2024)
Unemployment Monthly 32 cities + Cundinamarca (2021–2024)
Economically Active Population (PEA) Monthly 32 cities + Cundinamarca (2021–2024)
Working-Age Population (PET) Monthly 32 cities + Cundinamarca (2021–2024)
Total Population Monthly 32 cities + Cundinamarca (2021–2024)
Inactive Population Monthly 32 cities + Cundinamarca (2021–2024)
Participation / Employment / Unemployment / Inactivity Rates Monthly 32 cities + Cundinamarca (2021–2024)
Note: Author’s own construction.
@@
Consumer Price Index (IPC, Índice de Precios al Consumidor) Monthly National (1991-2025)
Producer Price Index (IPP, Índice de Precios al Productor) Monthly National (1991-2025)
Representative Market Exchange Rate (TRM, Tasa Representativa del Mercado) Daily / Monthly National (1991-2025)
Monthly Inflation (variation of IPC) Monthly National (1991-2025)
Note: Author’s own construction.
Appendix J.6 Official Informality Indicators
Table J.10: Labor Informality Indicators: National and Cities
Indicator Frequency Coverage
Informality Rate (Methodology I) Monthly National total (2007–2020)
Informality Rate (Methodology II) Monthly National total (2021–Present)
Informality Rate (Methodology I) Monthly 23 main cities (2007–2020)
Informality Rate (Methodology II) Monthly 23 main cities (2021–Present)
Note: Author’s own construction.